# Startup Acquisitions: Acquihires and Talent Hoarding

### v3

@@ 1 Introduction
Talent hoarding in our model is driven by the preemption effect, with firms acquiring talent partly to prevent competitors from becoming stronger. Many papers have identified the preemption effect in various settings, from patent races (Gilbert and Newbery, 1982), to M&As in the international-trade setting (Norbäck and Persson, 2004, 2007), to technology acquisitions (Bryan and Hovenkamp, 2020a), to the case of duopolists facing capacity constrained suppliers in a decentralized market (Burguet and Sákovics, 2017). Indeed, the point of our paper is to highlight that preemptive motives can be at work even when a startup operating in a different market is being acquired. However, preemptively acquiring labor is different from acquisitions of physical assets or technology because firms do not acquire property rights over labor. In our setting, this implies that some solutions to preemption, such as licensing (Katz and Shapiro, 1987), do not apply. Moreover, additional issues, such as the impact of the business cycle discussed in our thrid result, appear.
@@
Our second result concerns the effect of acquihires on consumer surplus. While increasing the efficiency of the acquirer, an acquihire leads to a loss of surplus induced by the disappearance of the startup. Assuming that part of the efficiency gain is passed on to consumers, the overall effect of the acquihire on consumer surplus is ambiguous. In particular, if high-match acquisitions increase it while low-match acquisitions decrease it, competition authorities who cannot identify match quality may face complex challenges in regulating acquihires. For instance, the welfare effects of banning acquihires might not depend straightforwardly on the ex-ante likelihood of a high-match deal. As we show, prohibiting acquihires tends to decrease consumer welfare when that probability is either very high or very low. In the former case, even though low-match firms have a strong incentive to hoard talent, the potential negative impact rarely materializes. In the latter case, low-match firms endogenously choose not to engage in (expensive) talent hoarding, so all observed acquihires are with high-match firms and thus welfare enhancing. It is when the probability of a high match is intermediate – such that low-match firms remain tempted to hoard talent and are not rare – that banning acquihires has the largest scope for enhancing expected consumer welfare.
@@
Our final result is that the labor-market outcomes (hiring, layoffs, and unemployment) for acquihired employees may become more volatile because of firms’ talent hoarding. To obtain this result, we expand our baseline model by adding a second period. In between periods, the economy may fall into a recession, and consequently, firms may get hit by potentially correlated adverse shocks. We show that relative to a benchmark case in which firms have no motive to hoard talent, talent hoarding always leads to more hiring and may also lead to more layoffs and unemployment for acquihired employees when the adverse shocks are sufficiently likely or sufficiently positively correlated. This finding lends support to the view that talent hoarding was a major contributing factor to the substantial number of layoffs in the tech industry following the increased downward pressure on the US economy in 2022.333See “Tech’s Talent Wars Have Come Back to Bite It,” New York Times, November 2022.
@@
We explore several extensions to our baseline model in the Online Appendix. These extensions include allowing the acquiring firm to differentiate between startup talent and technology, incorporating asymmetry in market power among potential acquirers, expanding the number of firms, introducing uncertainty in the order of moves, and accounting for partial acquisitions. Toward the paper’s end, we discuss these extensions in detail, arguing that they not only confirm the robustness of our core findings but provide additional insights. For instance, our extension with asymmetric firms highlights that, under reasonable profit assumptions, dominant firms have stronger incentives to hoard talent – an observation that resonates with concerns expressed by many regulators regarding the overhiring prevalent among tech giants.
@@
Our paper is most closely related to the literature studying the economics of startup acquisitions. Much of the early literature examined, in various settings, how the prospect of an acquisition affects the incentives of startups and incumbents to invest in innovation (e.g., Mason and Weeds, 2013; Norbäck and Persson, 2012; Gans and Stern, 2000; Phillips and Zhdanov, 2013; Rasmusen, 1988). Following Cunningham et al. (2021), who demonstrated that incumbents may acquire startups for anticompetitive reasons, a large literature has studied the effects a more restrictive merger policy would have on innovation and overall welfare (Motta and Peitz, 2021; Cabral, 2020, 2023; Katz, 2021; Letina et al., 2023). Others examine how acquisitions can steer the direction of innovation (Bryan and Hovenkamp, 2020a; Callander and Matouschek, 2022; Dijk et al., 2021), and Fumagalli et al. (2023) consider the impact of financial constraints. Several papers consider dynamic incentives (Cabral, 2018; Bryan and Hovenkamp, 2020a; Hollenbeck, 2020; Denicolo and Polo, 2021). A key insight is that if the incumbent pulls too far ahead in the technology space, the pace of innovation will go down. There is also the possibility that the incumbent creates a kill-zone that disincentivizes entry, either by acquiring entrants, copying their products, or heavily investing in innovation (Kamepalli et al., 2021; Teh et al., 2022; Shelegia and Motta, 2021; Bao and Eeckhout, 2023). On the empirical side, Ederer and Pellegrino (2023) show that startups increasingly favor acquisitions over IPOs as exit strategies. Finally, several papers empirically study acquisitions in the tech sector (Affeldt and Kesler, 2021a, b; Prado and Bauer, 2022; Jin et al., 2023; Gugler et al., 2023; Eisfeld, 2022; Gautier and Lamesch, 2021). Our paper differs from this literature by considering startups that are not potential competitors of the incumbents. The main channel through which acquisitions create inefficiencies is thus fundamentally different.
@@
Our paper also relates to the empirical literature that directly studies acquihires. Ouimet and Zarutskie (2020), Ng and Stuart (2021), Chen et al. (2021), and Chen et al. (2022) show that acquiring talent is indeed an important motivation for acquisitions. However, acquihired employees separate at a higher rate than regularly hired employees (Ng and Stuart, 2021; Verginer et al., 2022), possibly out of a preference for working at startups or a misalignment with the acquirer’s plans (Kim, 2020; Loh et al., 2019). This empirical finding is consistent with our theoretical result.
@@ 2 Model
Finally, the timing is as follows. At the outset, nature draws the private match types of the firms. In the first stage, firm 1 has the opportunity to attempt an acquihire. The entrepreneur can accept or reject the bid. If the entrepreneur accepts the bid, the game ends. Otherwise, we move to the second stage. In this stage, firm 2 has the opportunity to attempt an acquihire. The entrepreneur can accept or reject the bid, after which the game ends. [^6] In Online Appendix B.5 we show that the emergence of incentives to hoard talent does not hinge on firms’ knowledge of the order of moves nor on the fact that the firm gets the full surplus from the acquihire.
We expect our model to be more applicable to situations in which the startup is still relatively young and not widely known, so that firms becoming aware of it sequentially is appropriate. Should the startup already be more mature or the founder(s) actively looking for an exit, a simultaneous-move model may be more suitable. In such instances, talent hoarding, as outlined in Proposition 1 below, is less likely to occur, as a firm with high match value should be able to outbid a low-match-value competitor.
@@ 3 Talent Hoarding
The intuition for Proposition 2 (iii) is that when $CS_{H}>CS_{F}+CS_{E}>CS_{L}$, acquihires are harmful only when there is talent hoarding (i.e., both low- and high-match firms engage in an acquihire, requiring $\lambda\geq\lambda_{A}$) and the probability of a high match is sufficiently low (requiring $\lambda<\lambda_{CS}$). Hence, consumer-surplus-destroying acquihires can occur only for intermediate $\lambda$, that is, when $\lambda\in[\lambda_{A},\lambda_{CS})$.
Figure 1 illustrates Proposition 2(iii) using our Cournot example with $\pi_{E}=0.9$. The two panels only differ in the level of consumer surplus generated by the startup. We set $CS_{E}=0.4$ for the left panel and $CS_{E}=0.5$ for the right panel. In both panels, as $\lambda$ grows from 00 to $\lambda_{A}$, the consumer surplus when acquisitions are allowed increases (the solid line). For these parameters, firms endogenously only engage in high-match acquihires, which benefit consumers. At $\lambda_{A}$, low-match firms start talent hoarding, causing a discontinuous drop in consumer surplus visible on both panels. However, in the left panel $\lambda_{CS}<\lambda_{A}$, so the drop at $\lambda_{A}$ is not sufficient to lower the consumer surplus below the level achieved when acquisitions are prohibited (the dash-dotted line). In the right panel, $\lambda_{A}<\lambda_{CS}$, so that for all $\lambda\in[\lambda_{A},\lambda_{CS})$ the average consumer surplus is lower when acquihires are allowed than when they are not. As $\lambda$ increases beyond $\lambda_{CS}$, high-match acquihires become so likely that the overall consumer surplus is again higher than when acquisitions are prohibited. Thus, when $CS_{H}>CS_{F}+CS_{E}>CS_{L}$, allowing acquihires only lowers consumer surplus for intermediate values of $\lambda$. Finally, the dashed line represents the consumer surplus that could be achieved if the regulators could differentiate between the low-match and high-match acquihires. In that case, allowing only high-match acquihires always increases consumer surplus. Of course, if the regulators can only imperfectly detect match types, that lowers the expected consumer surplus below the dashed line.
@@
4 Hiring, Separation, and Unemployment
We now discuss the implications of talent hoarding on the hiring, separation, and unemployment of acquihired employees. To do so, we expand our baseline model by introducing a second period and allowing for economic downturns between the two periods.
The first period of this expanded model is identical to the baseline model in Section 2: the firms’ match types are private information, and firm 1 has the opportunity to pursue an acquihire before firm 2. Between the periods, the economy suffers a downturn with probability $\delta\in(0,1)$. If a downturn materializes – an event that is publicly observable – the firms may be hit by adverse shocks (see details below). In period 2, the entrepreneur has the option of creating a new startup, once more leading to an outside option of $\pi_{E}$ for her. [^10] There is empirical evidence that acquihired employees who leave the acquirer are likely to join a new startup (Ng and Stuart, 2021; Kim, 2022, 2020). If the entrepreneur was employed by a firm in period 1, that firm must decide whether to continue the relationship (at the cost of $\pi_{E}$) or lay off the entrepreneur, who might then be hired by the other firm. If there was no acquihire in period 1, it is again firm 1 that moves first in period 2.
The adverse shocks come with a commonly known joint probability distribution over “downgrades” for firms, though each firm only observes the realization of its own shock. More specifically, firm $i$ is hit by a shock $S_{i}\in\{D,N\}$, where it is either downgraded from high to low match (if possible) or not affected by the shock. Thus, if a low-match firm is hit by a downgrade, it stays a low-match firm, while a high-match firm turns into a low-match firm. If a firm is not affected by the shock, its match quality stays the same. Let $(S_{1},S_{2})\in\{D,N\}^{2}$ be the profile of shocks hitting the firms, which follows the distribution
@@ 4 Hiring, Separation, and Unemployment
To make a comparative statement, we need a benchmark relative to which we can compare the hiring, separation, and unemployment in our model with talent hoarding. To do so, consider the case in which $\Pi_{F}=\underline{\Pi}_{F}^{H}=\underline{\Pi}_{F}^{L}$ so that an acquihire by firm $i$ does not affect firm $j$’s profits. Thus, there are no incentives to hoard talent in this benchmark.
Before we state the formal result, consider the following intuition. If the entrepreneur was hired by a high-match firm in period 1 and this firm is not affected by the economic downturn (i.e., remains high-match), the firm will continue to employ the entrepreneur. Therefore, no separation or unemployment is observed in that case. [^11] In the video game industry, Loh et al. (2019) document that when the skills of the employees and the needs of the acquirer match well, the employees are more likely to stay with the acquirer. In contrast, if the entrepreneur was hired by a low-match firm in period 1 or a downgraded high-match firm, several period 2 outcomes can arise. Talent-hoarding motives may induce the continued employment of the entrepreneur if the competitor is believed to have a high match value with a high enough probability. Thus, we do not observe any separation or unemployment of the entrepreneur. Otherwise, we may observe a layoff of the entrepreneur, who is subsequently hired by a high-match competitor. Hence, while we observe separation, the entrepreneur does not become unemployed. Finally, the entrepreneur may be laid off and not hired by the competitor, resulting in both separation and unemployment. [^12] The term “unemployment” here means that the entrepreneur is laid off and then not employed by the competitor. Since the entrepreneur can start their own business, unemployment in the precise sense of the term does not occur. Note that these distinct period 2 outcomes in turn affect the behavior of low-match firms in period 1, changing the acquihire threshold. Solving the game fully and deriving the probabilities of period 1 hiring as well as observing separation and unemployment in period 2, we obtain the following result, proved in the appendix.
@@
The presence of talent-hoarding motives always leads to more hiring than in the benchmark. Additionally, provided that $\min\left\{\frac{\lambda_{A}}{\lambda},\frac{1-\lambda}{\lambda}\right\}>(1-r)(1-\gamma)$, talent hoarding also leads to more separation and unemployment than in the benchmark.

### v4

@@ 1 Introduction
Talent hoarding in our model is driven by the preemption effect, with firms acquiring talent partly to prevent competitors from becoming stronger. Many papers have identified the preemption effect in various settings, from patent races (Gilbert and Newbery, 1982), to M&As in the international-trade setting (Norbäck and Persson, 2004, 2007), to technology acquisitions (Bryan and Hovenkamp, 2020a), to the case of duopolists facing capacity constrained suppliers in a decentralized market (Burguet and Sákovics, 2017). Indeed, the point of our paper is to highlight that preemptive motives can be at work even when a startup operating in a different market is being acquired. However, preemptively acquiring labor is different from acquisitions of physical assets or technology because firms do not acquire property rights over labor. In our setting, this implies that some solutions to preemption, such as licensing (Katz and Shapiro, 1987), do not apply. Moreover, additional issues, such as the impact of the business cycle discussed in our third result, appear.
@@
Our second main result concerns the effect of acquihires on consumer surplus. While increasing the efficiency of the acquirer, an acquihire leads to a loss of surplus induced by the disappearance of the startup. Assuming that part of the efficiency gain is passed on to consumers, the overall effect of the acquihire on consumer surplus is ambiguous. In particular, if high-match acquisitions increase it while low-match acquisitions decrease it, competition authorities who cannot identify match quality may face complex challenges in regulating acquihires. For instance, the welfare effects of banning acquihires might not depend straightforwardly on the ex-ante likelihood of a high-match deal. As we show, prohibiting acquihires tends to decrease consumer welfare when that probability is either very high or very low. In the former case, even though low-match firms have a strong incentive to hoard talent, the potential negative impact rarely materializes. In the latter case, low-match firms endogenously choose not to engage in (expensive) talent hoarding, so all observed acquihires are with high-match firms and thus welfare enhancing. It is when the probability of a high match is intermediate – such that low-match firms remain tempted to hoard talent and are not rare – that banning acquihires has the largest scope for enhancing expected consumer welfare.
@@
Our final main result is that the labor-market outcomes (hiring and layoffs) for acquihired employees may become more volatile because of firms’ talent hoarding. To obtain this result, we expand our baseline model by adding a second period. In between periods, the economy may fall into a recession, and consequently, firms may get hit by potentially correlated adverse shocks. We show that relative to a benchmark case in which firms have no motive to hoard talent, talent hoarding always leads to more hiring and may also lead to more layoffs for acquihired employees when the adverse shocks are sufficiently likely or sufficiently positively correlated. This finding lends support to the view that talent hoarding was a major contributing factor to the substantial number of layoffs in the tech industry following the increased downward pressure on the US economy in 2022.333See “Tech’s Talent Wars Have Come Back to Bite It,” New York Times, November 2022.
Talent hoarding is likely to be a problem whenever firms can identify highly impactful individuals before their competitors do so. As such, it can, in principle, also materialize when firms hire new employees who possess a scarce and desirable skill set. In such situations, the insights of our model will also apply. However, we focus on acquihires (and not hiring in general) because talent hoarding is likely to be particularly acute in this setting, as it combines individuals who may not be known in the industry yet (the entrepreneurs) with a verifiable signal of their ability (the startup that they built).
@@
We explore two important extensions in Section 5. First, we consider the situation where a startup holds value for firms due to both its employees and its technology. This extension highlights that, unlike labor, technology can be sold or licensed following an acquisition since firms do not obtain property rights over labor. Second, we examine partial acquisitions or investments in startups. In these instances, the acquiring firm purchases only a portion of the startup, gaining a share of its profits and some decision-making authority. We demonstrate that when this decision power includes the ability to block acquisitions by competitors, such investments can be an attractive alternative to full acquihires. We also show that partial acquisitions generate lower inefficiencies than full acquihires but that they occur more often. Beyond these two extensions, we discuss in Section 6 how the emergence of talent hoarding is robust to incorporating asymmetry in market power among potential acquirers, expanding the number of firms, introducing uncertainty in the order of moves, and allowing for some surplus sharing between the acquiring firm and the entrepreneur. Overall, these extensions not only confirm the robustness of our core findings but provide additional insights. For instance, our extension with asymmetric firms highlights that, under reasonable profit assumptions, dominant firms have stronger incentives to hoard talent – an observation that resonates with concerns expressed by many regulators regarding the excessive hiring prevalent among tech giants.
@@
Our paper is most closely related to the literature studying the economics of startup acquisitions. Much of the early literature examined, in various settings, how the prospect of an acquisition affects the incentives of startups and incumbents to invest in innovation (e.g., Mason and Weeds, 2013; Norbäck and Persson, 2012; Gans and Stern, 2000; Phillips and Zhdanov, 2013; Rasmusen, 1988). Following Cunningham et al. (2021), who demonstrated that incumbents may acquire startups for anticompetitive reasons, a large literature has studied the effects a more restrictive merger policy would have on innovation and overall welfare (Motta and Peitz, 2021; Cabral, 2020, 2023; Katz, 2021; Letina et al., 2024). Others examine how acquisitions can steer the direction of innovation (Bryan and Hovenkamp, 2020a; Callander and Matouschek, 2022; Dijk et al., 2024), and Fumagalli et al. (2023) consider the impact of financial constraints. Several papers consider dynamic incentives (Cabral, 2018; Bryan and Hovenkamp, 2020a; Hollenbeck, 2020; Denicolo and Polo, 2021). A key insight is that if the incumbent pulls too far ahead in the technology space, the pace of innovation will go down. There is also the possibility that the incumbent creates a kill-zone that disincentivizes entry, either by acquiring entrants, copying their products, or heavily investing in innovation (Kamepalli et al., 2021; Teh et al., 2022; Shelegia and Motta, 2021; Bao and Eeckhout, 2023). On the empirical side, Ederer and Pellegrino (2023) show that startups increasingly favor acquisitions over IPOs as exit strategies. Finally, several papers empirically study acquisitions in the tech sector (Affeldt and Kesler, 2021a, b; Prado and Bauer, 2022; Jin et al., 2023; Gugler et al., 2023; Eisfeld, 2022; Gautier and Lamesch, 2021). Our paper differs from this literature by considering startups that are not potential competitors of the incumbents. The main channel through which acquisitions create inefficiencies is thus fundamentally different.
@@
Our paper also relates to the empirical literature that directly studies acquihires. Ouimet and Zarutskie (2020), Ng and Stuart (2021), Chen et al. (2021), and Chen et al. (2023) show that acquiring talent is indeed an important motivation for acquisitions. However, acquihired employees separate at a higher rate than regularly hired employees (Ng and Stuart, 2021; Verginer et al., 2022), possibly out of a preference for working at startups or a misalignment with the acquirer’s plans (Kim, 2023; Loh et al., 2019). This empirical finding is consistent with our theoretical result.
@@ 2 Model
Finally, the timing is as follows. At the outset, nature draws the private match types of the firms. In the first stage, firm 1 has the opportunity to attempt an acquihire. The entrepreneur can accept or reject the bid. If the entrepreneur accepts the bid, the game ends. Otherwise, we move to the second stage. In this stage, firm 2 has the opportunity to attempt an acquihire. The entrepreneur can accept or reject the bid, after which the game ends. [^6] We discuss in Section 6 that the emergence of incentives to hoard talent does not hinge on firms’ knowledge of the order of moves nor on the fact that the firm gets the full surplus from the acquihire.
We expect our model to be more applicable to situations in which the startup is still relatively young and not widely known, so that firms becoming aware of it sequentially is appropriate. Indeed, as evidenced in the report issued by the Federal Trade Commission’s report on non-Hart-Scott-Rodino acquisitions by Alphabet, Amazon.com, Apple, (then still) Facebook, and Microsoft from 2010-2019, such situations seem to be quite prevalent if not the norm (FTC, 2021). The report shows that (i) approximately 65% of all transactions cost less than \$25m, (ii) almost half of all transactions comprised startups that were less than five years old, (iii) more than 50% of the transactions involved targets with fewer than ten employees. Overall, this supports the notion of small, relatively young and not widely known startups. [^7] Should the startup already be more mature or the founder(s) actively looking for an exit, a simultaneous-move model may be more suitable. In such instances, talent hoarding, as outlined in Proposition 1 below, is less likely to occur, as a firm with high match value should be able to outbid a low-match-value competitor.
@@ 3 Talent Hoarding
The intuition for Proposition 2 (iii) is that when $CS_{H}>CS_{F}+CS_{E}>CS_{L}$, acquihires are harmful only when there is talent hoarding (i.e., both low- and high-match firms engage in an acquihire, requiring $\lambda\geq\lambda_{A}$) and the probability of a high match is sufficiently low (requiring $\lambda<\lambda_{CS}$). Hence, acquihires lower expected consumer surplus only for intermediate $\lambda$, that is, when $\lambda\in[\lambda_{A},\lambda_{CS})$.
Figure 1 illustrates Proposition 2(iii) using our Cournot example with $\pi_{E}=0.9$. The two panels only differ in the level of consumer surplus generated by the startup. We set $CS_{E}=0.4$ for the left panel and $CS_{E}=0.5$ for the right panel. In both panels, as $\lambda$ grows from 00 to $\lambda_{A}$, the consumer surplus when acquisitions are allowed increases (the solid line). For these parameters, firms endogenously only engage in high-match acquihires, which benefit consumers. At $\lambda_{A}$, low-match firms start talent hoarding, causing a discontinuous drop in consumer surplus visible on both panels. However, in the left panel, $\lambda_{CS}<\lambda_{A}$, so the drop at $\lambda_{A}$ is not sufficient to lower the consumer surplus below the level achieved when acquisitions are prohibited (the dash-dotted line). In the right panel, $\lambda_{A}<\lambda_{CS}$, so that for all $\lambda\in[\lambda_{A},\lambda_{CS})$ the average consumer surplus is lower when acquihires are allowed than when they are not. As $\lambda$ increases beyond $\lambda_{CS}$, high-match acquihires become so likely that the overall consumer surplus is again higher than when acquisitions are prohibited. Thus, when $CS_{H}>CS_{F}+CS_{E}>CS_{L}$, allowing acquihires only lowers consumer surplus for intermediate values of $\lambda$. Finally, the dashed line represents the consumer surplus that could be achieved if the regulators could differentiate between the low-match and high-match acquihires. In that case, allowing only high-match acquihires always increases consumer surplus. Of course, if the regulators can only imperfectly detect match types, that lowers the expected consumer surplus below the dashed line.
@@
4 Hiring and Layoffs
We now discuss the implications of talent hoarding on the hiring and layoffs of acquihired employees. To do so, we expand our baseline model by introducing a second period and allowing for economic downturns between the two periods.
The first period of this expanded model is identical to the baseline model in Section 2: the firms’ match types are private information, and firm 1 has the opportunity to pursue an acquihire before firm 2. Between the periods, the economy suffers a downturn with probability $\delta\in(0,1)$. If a downturn materializes – an event that is publicly observable – the firms may be hit by adverse shocks (see details below). In period 2, the entrepreneur has the option of creating a new startup, once more leading to an outside option of $\pi_{E}$ for her. [^11] There is empirical evidence that acquihired employees who leave the acquirer are likely to join a new startup (Ng and Stuart, 2021; Kim, 2022, 2023). If the entrepreneur was employed by a firm in period 1, that firm must decide whether to continue the relationship (at the cost of $\pi_{E}$) or lay off the entrepreneur, who might then be hired by the other firm. If there was no acquihire in period 1, it is again firm 1 that moves first in period 2.
The adverse shocks come with a commonly known joint probability distribution over “downgrades” for firms, though each firm only observes the realization of its own shock. More specifically, firm $i$ is hit by a shock $S_{i}\in\{D,N\}$, where it is either downgraded from high to low match (if possible) or not affected by the shock. Thus, if a low-match firm is hit by a downgrade, it stays a low-match firm, while a high-match firm turns into a low-match firm. If a firm is not affected by the shock, its match quality stays the same. [^12] Of course, an economy-wide adverse shock may affect more parameters than just the profitability of the high-match firm (for example, it may impact the profitability of startups or firms even in the absence of any acquisitions). We are focusing on the impact on the high-match firms as they will presumably seek to expand, the profitability of which is particularly likely to be impacted by a negative shock to the economy. Let $(S_{1},S_{2})\in\{D,N\}^{2}$ be the profile of shocks hitting the firms, which follows the distribution
@@ 4 Hiring and Layoffs
To make a comparative statement, we need a benchmark relative to which we can compare the hiring and layoffs in our model with talent hoarding. To do so, consider the case in which $\Pi_{F}=\underline{\Pi}_{F}^{H}=\underline{\Pi}_{F}^{L}$ so that an acquihire by firm $i$ does not affect firm $j$’s profits. Thus, there are no incentives to hoard talent in this benchmark.
Before we state the formal result, consider the following intuition. If the entrepreneur was hired by a high-match firm in period 1 and this firm is not affected by the economic downturn (i.e., remains high-match), the firm will continue to employ the entrepreneur. Therefore, no layoff is observed in that case. [^13] In the video game industry, Loh et al. (2019) document that when the skills of the employees and the needs of the acquirer match well, the employees are more likely to stay with the acquirer. In contrast, if the entrepreneur was hired by a low-match firm in period 1 or a downgraded high-match firm, several period 2 outcomes can arise. Talent-hoarding motives may induce the continued employment of the entrepreneur if the competitor is believed to have a high match value with a high enough probability. Thus, we do not observe any layoff of the entrepreneur. Otherwise, we may observe a layoff of the entrepreneur, who is subsequently hired by a high-match competitor. Hence, while we observe a separation, the entrepreneur remains employed in the industry. Finally, the entrepreneur may be laid off and not hired by the competitor, resulting in the entrepreneur leaving the industry. Note that these distinct period 2 outcomes, in turn, affect the behavior of low-match firms in period 1, changing the acquihire threshold. Solving the game fully and deriving the probabilities of period 1 hiring as well as observing a layoff in period 2, we obtain the following result, proved in the appendix.
@@
The presence of talent-hoarding motives always leads to more hiring than in the benchmark. Additionally, provided that $\min\left\{\frac{\lambda_{A}}{\lambda},\frac{1-\lambda}{\lambda}\right\}>(1-r)(1-\gamma)$, talent hoarding also leads to more layoffs than in the benchmark.